# Yiqi Lin1* Conghui He1*† Alex Jinpeng Wang2* Bin Wang1* Weijia Li3 Mike Zheng Shou2

### 28 Dec 2023 v2

@@ Abstract
Despite CLIP being the foundation model in numerous vision-language applications, the CLIP suffers from a severe text spotting bias. Such bias causes CLIP models to ‘Parrot’ the visual text embedded within images while disregarding the authentic visual semantics. We uncover that in the most popular image-text dataset LAION-2B, the captions also densely parrot (spell) the text embedded in images. Our analysis shows that around 50% of images are embedded with visual text content, and 90% of their captions more or less parrot the visual text. Based on such observation, we thoroughly inspect the different released versions of CLIP models and verify that the visual text is the dominant factor in measuring the LAION-style image-text similarity for these models. To examine whether these parrot captions shape the text spotting bias, we train a series of CLIP models with LAION subsets curated by different parrot-caption-oriented criteria. We show that training with parrot captions easily shapes such bias but harms the expected visual-language representation learning in CLIP models. This suggests that it is urgent to revisit either the design of CLIP-like models or the existing image-text dataset curation pipeline built on CLIP score filtering.
@@ 3 Terminology
  • Text Removal Image: do inpainting in the specific spotted text area (All-Emb., Co-Emb., or Random). The random is implemented by sampling other image’s text areas. For the different inpainting results, see Fig. 2
@@
Number of Total Images 1,985,284,122
Number of Images w/ Emb. Text 1,083,896,427
Image w/ Emb. Text Rate 54.60%
Co-Emb. Text Rate (in Total) 15.42%
– (in Images w/ Emb. Text) 28.24%
Fuzzy Co-Emb. Text Rate (in Total) 30.46%
– (in Images w/ Emb. Text) 55.79%
@@ 4.1 Implementation Details
Text Spotting and Matching: To detect and recognize text across various scenes, we adopt DeepSolo [37] as our text-spotting model and use the pre-trained checkpoints with the ViTAEv2-S [40] backbone in default setting. The output format of the text spotting model is a sequence of polygons of text location and their recognized characters. To identify the correlation between the spotted text and captions, we use Algorithm 1 to calculate the Co-Emb. text rate in each image-text pair. Considering the predictions that the text spotting model might miss or misspell words in complex scenes, we also use Levenshtein distance to calculate the fuzzing similarity and reported in Tab. 1.
@@ 4.2 Statistic and Observations from LAION-2B
The overall statistics of the 2 billion image-text pairs are reported in Tab. 1. In summary, the images embedded with visual text content reach a surprisingly high proportion of 54.60% in the investigated data. Around 15% of words in the dataset captions are Co-Emb. text, and the proportion of Co-Emb. text can further reach 30% when considering the fuzzy matching results of the spotted text and captions. This suggests that the CLIP models trained on these data might lead to a high bias toward text spotting. To better visualize the data distribution, we provide cluster-specific statics results and top CLIP score samples of text-dominated clusters in Fig. 3. We divide all images into 100 clusters based on visual similarity and classify them into three types of data according to the OCR results. Every cluster contains more or less images embedded with text. Combined with sample visualization, we observe that in the LAION collected data, the parrot captions cover various scenes. In the subsets of images embedded with text, around 90% of captions at least parrot one concurrent word appearing in the image. It suggests that the data collection pipeline of LAION [28] has a strong bias to introduce parrot captions from web data.
@@ 5 Inspecting Pre-Trained CLIP Models
(a) Statistic of 1-gram vocabularies.
(b) Grouped by 1-gram frequency intervals.
(c) Grouped by Caption 1-gram length.
(d) Grouped by Caption 2-gram length.
Figure 5: OpenCLIP more bias towards text spotting than the CLIP model. Grouped score distributions of prompting CLIP and OpenCLIP models with N-gram Syn-Emb. text and synthetic images for model preference investigation.
@@ 5.1 Ablation of Embedded Text Removal
Text Removal via Inpainting: Given the OCR predicted results, we use the fast marching method [31] to do text in-painting in the area of the spotted text polygons. Accordingly, we generate two versions of text removal results for each image with embedded text, i.e., All-Emb. text removal and Co-Emb. text removal, as the parrot caption prediction is imperfect due to the limitation of OCR models. We also generate random inpainting images with randomly sampled spotted text polygons from other images to ablate the information loss caused by image inpainting. The spotted text masks and inpainting results are shown in Fig 2.
@@
Results: Firstly, we examine whether the CLIP models prefer recognizing more commonly seen words (with high frequency in vocabulary). Therefore, we group the 1-gram results based on their frequency interval in the whole vocabulary, as shown in Fig. 4(b). The results show that the OpenCLIP model clearly has a stronger text spotting capacity than CLIP, i.e., more biased towards text spotting. We also observe that all the CLIP models are more sensitive to the vocabulary built from the concurrent words. Interestingly, both CLIP and OpenCLIP models have slightly higher scores on the less frequent grams. Secondly, considering the long-tail grams might contain more characters, we further group the 1-gram and 2-gram results based on their text length in Fig. 4(c) and Fig. 4(d). Note that the Co-Emb. text is not regularly arranged in the images, making it hard to extract continuous word sequences. Results show that all the models are better at spotting the longer words, possibly due to the tokenizer used in the text encoder, making them more discriminative. Meanwhile, in the groups of 2-gram samples, the scores gradually drop when spotting the extremely long text, indicating that the spotting capacity of CLIP models is possibly built on word-by-word.
@@
6 Training CLIP on Emb. Text Curated Data
6.1 Experiment Setups
@@
Evaluation: We follow the DataComp benchmark [8] using 38 zero-shot classification and retrieval tasks as evaluation. We report the average performance (Avg.) of the DataComp benchmark and two subset track performances, ImageNet (IN) and Retrieval (Ret.). To evaluate the text spotting capacity, we use a synthetic benchmark the same as in Sec. 5.2 and a real-world benchmark sampled from LAION-2B as the validation set. In the synthetic benchmark, we calculate the similarity of all the 1-gram synthetic image-text pairs from caption vocabulary and report all the trained model results in Fig 6. For the real-world benchmark, we sample 1M image-text pairs without any embedded text and 1M samples dominated by the parrot caption (the relative scores between raw and Co-Emb. text removal images higher than 0.2). Fig. 7 aggregates the mean scores of the 2M subset and reports the mean scores with text removal results.
6.2 Ablation Study
@@
Curation II: Co-Emb. Text Rate (CoTR). Tab. 2 reports the CLIP models trained on parrot captions with different CoTR. We first select all the images with embedded text and then sample images based on the CoTR depicted at Algorithm 1 with different thresholds. With increasing CoTR, all the zero-shot benchmark performance drops significantly. Despite the images in the subset (CoTR = 0) all embedded with text, the pre-trained model performs similarly to the random baseline in 2. It indicates that the parrot caption is more crucial than embedded text in reducing the pre-trained data quality. For the text spotting capacity, Fig. 6 and 7 show that the increasing CoTR does not lead to stronger text spotting capacity, possibly due to the average length of captions decreasing in higher CoTR data.
Number of Total Images 516166764
Number of Images w/ Emb. Text 259101894
@@
1-gram Sync. Score (mean) $\downarrow$ 0.163 0.317 0.368 0.268 0.338
1-gram Sync. Score (std) 0.659 0.305 0.427 0.247 0.341
ImageNet 0.526 0.633 0.666 0.176 0.459
ImageNet dist. shifts 0.404 0.485 0.522 0.152 0.378
@@
Curation III: Relative Score from Text Removal. Given the observations in Sec. 5.1, we further select a series of subsets based on the relative score of images before and after text removal. The subsets with higher relative scores are more dominant embedded text (RSA) or parrot captions (RSC) in CLIP score measurement. In Tab. 4 and 5, we report the zero-shot performance of models trained on subsets with different relative score thresholds. The CLIP pre-trained with higher RSA or RSC both get worse downstream performance. Importantly, the average raw CLIP scores S($\bullet$) of these subsets have a positive correlation with RSA or RSC, indicating using CLIP scores from a biased pre-trained model as the data filtering strategy can be unreliable. When comparing the RSA and RSC, the results show that the samples dominated by the latter, i.e., parrot captions, are less informative for CLIP training. Moreover, Fig. 6 and. 7 show that the text spotting capacity of CLIP can be further improved by training on the samples using relative scores as data curation criteria against CoTR.
7 Profiling More Image-Text Dataset
@@ 13 Curation Ablation on 12M scales
We further provide the ablation study on 12M scales in Tab. 1112 and 13. All the results are consistent with the 3M scale results. Fig. 9 reports the text spotting capacity of models on 12M scales using the synthetic benchmark the same as Sec. 6.2. It shows that training with more parrot caption samples does not lead to a stronger text spotting performance in synthetic benchmarks.
@@ 14 Full Tracks Evaluation on DataComp
In Fig. 11,  12, and  13 we report all dataset results on DataComp [8] of the ablation study. In most vision-centric datasets, the model performance is consistent with the average performance. Meanwhile, the results also indicate that the model with stronger text capacity achieves better performance on the text-oriented task, such as MINST.

### 01 Feb 2024 v3

@@ Abstract
Despite CLIP being the foundation model in numerous vision-language applications, the CLIP suffers from a severe text spotting bias. Such bias causes CLIP models to ‘Parrot’ the visual text embedded within images while disregarding the authentic visual semantics. We uncover that in the most popular image-text dataset LAION-2B, the captions also densely parrot (spell) the text embedded in images. Our analysis shows that around 50% of images are embedded with visual text content, and around 30% of captions words are in these embedded visual content. Based on such observation, we thoroughly inspect the different released versions of CLIP models and verify that the visual text is the dominant factor in measuring the LAION-style image-text similarity for these models. To examine whether these parrot captions shape the text spotting bias, we train a series of CLIP models with LAION subsets curated by different parrot-caption-oriented criteria. We show that training with parrot captions easily shapes such bias but harms the expected visual-language representation learning in CLIP models. This suggests that it is urgent to revisit either the design of CLIP-like models or the existing image-text dataset curation pipeline built on CLIP score filtering.
@@ 3 Terminology
  • Text Removal Image: do inpainting in the specific spotted text area (All-Emb., Co-Emb., or Random). The random is implemented by sampling other image’s text areas. For the different inpainting results, see Fig. 2.
@@ 4.1 Implementation Details
Text Spotting and Matching: To detect and recognize text across various scenes, we adopt DeepSolo [44] as our text-spotting model and use the pre-trained checkpoints with the ViTAEv2-S [47] backbone in default setting. The output format of the text spotting model is a sequence of polygons of text location and their recognized characters. Despite its strong performance, we empirically find that DeepSolo can not handle the crowd scenes well (with more than 100 separate words) but is only a small proportion of the dataset ($\sim$2%). To identify the correlation between the spotted text and captions, we use Algorithm 1 to calculate the Co-Emb. text rate in each image-text pair. Considering the predictions that the text spotting model might miss or misspell words in complex scenes, we also use Levenshtein distance to calculate the fuzzing similarity and reported in Tab. 1.
Num. of Total Img. 1,985,284,122
Num. of Img. w/ Emb. Text 1,083,896,427
Num. of Img. w/ Co-Emb. Text 663,600,432
Co-Emb. Text Rate (in Total) 15.42%
– (in Img. w/ Emb. Text) 28.24%
Fuzzy Co-Emb. Text Rate (in Total) 30.46%
– (in Img. w/ Emb. Text) 55.79%
@@ 4.2 Statistic and Observations from LAION-2B
The overall statistics of the 2 billion image-text pairs are reported in Tab. 1. In summary, the images embedded with visual text content reach a surprisingly high proportion of 54.60% in the investigated data. Around 15% of words in the dataset captions are Co-Emb. text, and the proportion of Co-Emb. text can further reach 30% when considering the fuzzy matching results of the spotted text and captions. This suggests that the CLIP models trained on these data might lead to a high bias toward text spotting.
To better visualize the data distribution, we provide cluster-specific statics results and top CLIP score samples of text-dominated clusters in Fig. 3. We divide all images into 100 clusters based on visual similarity and classify them into three data types according to the OCR results. Every cluster contains more or less images embedded with text. Combined with sample visualization, we observe that in the LAION collected data, the parrot captions cover various scenes. In the subsets of images embedded with text, around 60% of captions at least precisely parrot one concurrent word (Co-Emb. Text Rate $>0$) appearing in the image. It suggests that the data collection pipeline of LAION [33] has a strong bias to introduce parrot captions from web data.
To better understand Co-Emb. Text, we provide a more thorough analysis of the word counting and text size of parrot captions. As shown in Fig. 4a, the results show that a large proportion of the Co-Emb. Text only takes a few words. However, we also find a large number of captions that are almost full parrot captions (see areas around the heatmap diagonal). Next, in Fig. 4b, we investigate the correlation between the size of concurrent words and CLIP score. The results show that the large text size does not usually lead to a higher score; meanwhile, the small text size can also dominate the score. One possible reason is the text content and input resolution may matter more for CLIP. Moreover, we discover that the larger text is more likely to be parroted in the captions, as shown in Fig. 4c.
@@ 5 Inspecting Pre-Trained CLIP Models
Figure 4: (a): Visualization of the number of caption words and associated spotted concurrent words based on precise word matching. (b): Distribution of total area of concurrent words placed in the image and its ViT-B CLIP score. (c): Distribution of text size of the single concurrent word and other spotted word.
@@ 5.1 Ablation of Embedded Text Removal
Text Removal via Inpainting: Given the OCR predicted results, we use the fast marching method [38] to do text in-painting in the area of the spotted text polygons. Accordingly, we generate two versions of text removal results for each image with embedded text, i.e., All-Emb. text removal and Co-Emb. text removal, as the parrot caption prediction is imperfect due to the limitation of OCR models. We also generate random inpainting images with randomly sampled spotted text polygons from other images to ablate the information loss caused by image inpainting. The spotted text masks and inpainting results are shown in Fig. 2.
@@
Setup CLIP Score
$\bullet$ Raw w/o Emb. Text 0.3223 $\pm$ 0.0078
$\bullet$ Raw w/ Emb. Text 0.3358 $\pm$ 0.0094
$\boldsymbol{\times}$ Random All-Emb. Text Removal 0.3260 $\pm$ 0.0057
$\boldsymbol{\times}$ All-Emb. Text Removal 0.2974 $\pm$ 0.0197
$\blacksquare$ Random Co-Emb. Text Removal 0.3341 $\pm$ 0.0051
$\blacksquare$ Co-Emb. Text Removal 0.2993 $\pm$ 0.0146
Table 2: Mean CLIP score of different setups of text removal.
@@
Discussion: Due to the text removal, the image distribution may shift from the CLIP training set. Therefore, we provide two random removal baselines to examine the effect of distribution shift. In Tab. 2, we report the mean scores of different setups. Results show that the random baselines are very close to the raw image baseline, indicating that the CLIP model is robust to the distribution shift caused by information loss in inpainted regions.
(a) Statistic of 1-gram vocabularies.
(b) Grouped by 1-gram frequency intervals.
Figure 6: OpenCLIP more bias than the CLIP model. Grouped score distributions of prompting CLIP and OpenCLIP models with N-gram Syn-Emb. text and synthetic images.
@@
Results: Firstly, we examine whether the CLIP models prefer recognizing more commonly seen words (with high frequency in vocabulary). Therefore, we group the 1-gram results based on their frequency interval in the whole vocabulary, as shown in Fig. 5(b). The results show that the OpenCLIP model clearly has a stronger text spotting capacity than CLIP, i.e., more biased towards text spotting. We also observe that all the CLIP models are more sensitive to the vocabulary built from the concurrent words. Interestingly, both CLIP and OpenCLIP models have slightly higher scores on the less frequent grams. Secondly, considering the long-tail grams might contain more characters, we further group the 1-gram and 2-gram results based on their text length in Fig. 6(a) and Fig. 6(b). Note that the Co-Emb. text is not regularly arranged in the images, making it hard to extract continuous word sequences. Results show that all the models are better at spotting the longer words, possibly due to the tokenizer used in the text encoder, making them more discriminative. Meanwhile, in the groups of 2-gram samples, the scores gradually drop when spotting the extremely long text, indicating that the spotting capacity of CLIP models is possibly built on word-by-word.
(a) Grouped by Caption 1-gram length.
(b) Grouped by Caption 2-gram length.
Figure 7: CLIPs are better at spotting the longer words. Score distributions of N-gram Syn-Emb. text grouped by text length.
6 Training CLIP on Emb. Text Curated Data
6.1 Experiment Setups
@@
Evaluation: We follow the DataComp benchmark [10] using 38 zero-shot classification and retrieval tasks as evaluation. We report the average performance (Avg.) of the DataComp benchmark and two subset track performances, ImageNet (IN) and Retrieval (Ret.). To evaluate the text spotting capacity, we use a synthetic benchmark the same as in Sec. 5.2 and a real-world benchmark sampled from LAION-2B as the validation set. In the synthetic benchmark, we calculate the similarity of all the 1-gram synthetic image-text pairs from caption vocabulary and report all the trained model results in Fig 8. For the real-world benchmark, we sample 1M image-text pairs without any embedded text and 1M samples dominated by the parrot caption (the relative scores between raw and Co-Emb. text removal images higher than 0.2). Fig. 9 aggregates the mean scores of the 2M evaluation set and also reports the mean scores of applying text removal on the 2M evaluation set results.
6.2 Ablation Study on Data Curation
@@
Visual Question Answering (Acc) Image Captioning (CIDEr) Text-to-Image Retrieval(R@1) Image-to-Text Retrieval(R@1)
BLIP Data (3M) VQAv2 TextVQA ST-VQA COCO TextCaps COCO TextCaps COCO TextCaps
Rand 71.07 15.36 10.48 115.6 53.7 48.91 56.34 65.46 72.45
w/ Emb. Text 68.94 19.05 12.65 108.9 92.1 42.89 70.1 58.5 81.42
w/o Emb. Text 71.22 13.65 9.29 116.2 41.5 49.96 31.83 66.5 48.7
CoTR = 0.0 71.11 13.97 9.75 116.3 44.6 49.55 38.05 66.08 54.57
CoTR $\geq$ 0.3 67.4 19.28 11.81 104.9 96.9 37.78 67.28 51.98 78.2
CoTR $\geq$ 0.5 67.02 19.64 12.38 102.7 94.1 35.94 65.24 50.32 76.94
CoTR $\geq$ 0.8 66.38 18.50 12.00 100.9 91.6 34.13 62.65 46.9 73.56
CoTR = 1.0 66.18 18.47 12.80 101.2 91.3 33.55 61.83 46.62 73.05
RSA $<$ 0.0 70.79 14.16 9.64 115.7 44.9 48.25 36.85 64.72 54.7
RSA $\geq$ 0.0 70.03 18.76 11.81 111.9 84.5 46.25 68.61 62.92 81.23
RSA $\geq$ 0.1 68.14 19.48 13.33 105.6 96.1 39.96 68.13 54.64 79.37
RSA $\geq$ 0.2 66.01 21.06 11.85 98.7 94.4 33.03 64.17 47.12 75.33
RSA $\geq$ 0.3 64.20 18.44 12.04 95.26 91.1 26.64 60.11 37.3 70.24
RSC $<$ 0.0 70.13 15.19 10.74 112.2 46.7 46.8 41.95 63.24 58.05
RSC $\geq$ 0.0 68.86 20.12 13.75 107.8 93.5 42.0 69.78 57.42 80.92
RSC $\geq$ 0.1 67.35 20.54 12.84 103.4 96.9 36.4 66.69 51.02 77.79
RSC $\geq$ 0.2 62.62 20.32 13.14 98.7 92.8 30.08 61.38 42.96 71.98
RSC $\geq$ 0.3 63.75 18.94 13.03 92.9 88.7 24.23 58.35 34.72 68.95
Table 7: BLIP downstream tasks performance of pre-training on different curated 3M subsets. The gray color represents tasks requiring the model to read the text from images.
Curation II: Co-Emb. Text Rate (CoTR). Tab. 3 reports the CLIP models trained on parrot captions with different CoTR. We first select all the images with embedded text and then sample images based on the CoTR depicted at Algorithm 1 with different thresholds. With increasing CoTR, all the zero-shot benchmark performance drops significantly. Despite the images in the subset (CoTR = 0) all embedded with text, the pre-trained model performs similarly to the random baseline in 3. It indicates that the parrot caption is more crucial than embedded text in reducing the pre-trained data quality. For the text spotting capacity, Fig. 8 and 9 show that the increasing CoTR does not lead to stronger text spotting capacity, possibly due to the average length of captions decreasing in higher CoTR data.
Curation III: Relative Score from Text Removal. Given the observations in Sec. 5.1, we further select a series of subsets based on the relative score of images before and after text removal. The subsets with higher relative scores are more dominant embedded text (RSA) or parrot captions (RSC) in CLIP score measurement. In Tab. 5 and 6, we report the zero-shot performance of models trained on subsets with different relative score thresholds. The CLIP pre-trained with higher RSA or RSC both get worse downstream performance. Importantly, the average raw CLIP scores S($\bullet$) of these subsets have a positive correlation with RSA or RSC, indicating using CLIP scores from a biased pre-trained model as the data filtering strategy can be unreliable. When comparing the RSA and RSC, the results show that the samples dominated by the latter, i.e., parrot captions, are less informative for CLIP training. Moreover, Fig. 8 and. 9 show that the text spotting capacity of CLIP can be further improved by training on the samples using relative scores as data curation criteria against CoTR.
6.3 Ablation Study on Text-Oriented Tasks
Inspired by [11], we further investigate the model behavior on downstream tasks requiring reading text, including Visual Question Answering (VQA), Image Captioning, and Text-Image Retrieval. Specifically, for the text-oriented tasks, we use Text VQA [37] and ST-VQA [3] for VQA, and TextCaps [36] for captioning and retrieval. Moreover, we also provide the same tasks on the datasets that only require the model to see, i.e., the natural image dataset. Similarly, we use VQAv2 [13] for VQA and COCO [7] for captioning and retrieval. We choose BLIP [22] for the ablation study instead of CLIP as it can be directly applied to all these tasks. We first pre-train the BLIP on different subsets with 10 epochs and then finetune 10 epochs for VQA, and 5 epochs for captioning and retrieval. As shown in Tab.7, training BLIPs to spot text can boost their performance on the downstream tasks requiring the model to read but impede the performance of downstream tasks only requiring the model to see, which are consistent with the observation on classification tasks. Nevertheless, when BLIPs mainly focus on reading, e.g. (RSA $\geq$ 0.3), their text-oriented and natural downstream performance also decreases. In other words, the parrot captions can benefit the text-orient downstream tasks while requiring careful data mixing trade-off.
7 Profiling More Image-Text Dataset
Number of Total Images 527156206
Number of Images w/ Emb. Text 264618807
@@ 7 Profiling More Image-Text Dataset
Sync. Score $\downarrow$ 0.163 $\pm$ 0.065 0.317 $\pm$ 0.030 0.368 $\pm$ 0.042 0.268 $\pm$ 0.024 0.338 $\pm$ 0.034
IN 0.526 0.633 0.666 0.176 0.459
IN dist. shifts 0.404 0.485 0.522 0.152 0.378
@@ 13 Curation Ablation on 12M scales
We further provide the ablation study on 12M scales in Tab. 1314 and 15. All the results are consistent with the 3M scale results. Fig. 11 reports the text spotting capacity of models on 12M scales using the synthetic benchmark the same as Sec. 6.2. It shows that training with more parrot caption samples does not lead to a stronger text spotting performance in synthetic benchmarks.
@@ 14 Full Tracks Evaluation on DataComp
In Fig. 13,  14, and  15 we report all dataset results on DataComp [10] of the ablation study. In most vision-centric datasets, the model performance is consistent with the average performance. Meanwhile, the results also indicate that the model with stronger text capacity achieves better performance on the text-oriented task, such as MINST.